### v1

# DCT-MARL: A Dynamic Communication Topology Based MARL Algorithm for Platoon Control

Abstract
With the rapid advancement of vehicular communication and autonomous driving technologies, connected vehicle platoon has emerged as a promising approach to improve traffic efficiency and driving safety. Reliable Vehicle-to-Vehicle (V2V) communication is critical to achieving efficient cooperative control. However, in real-world traffic environments, V2V links may suffer from time-varying delay and packet loss, leading to degraded control performance and even safety risks. To mitigate the adverse effects of non-ideal communication, this paper proposes a Dynamic Communication Topology based Multi-Agent Reinforcement Learning (DCT-MARL) algorithm for robust cooperative platoon control. Specifically, the state space is augmented with historical control action and delay to enhance robustness against communication delay. To mitigate the impact of packet loss, a multi-key gated communication mechanism is introduced, which dynamically adjusts the communication topology based on the correlation between agents and their current communication status. Simulation results demonstrate that the proposed DCT-MARL significantly outperforms state-of-the-art methods in terms of string stability and driving comfort, validating its superior robustness and effectiveness.
Index Terms:
Connected Vehicle Platoon, Dynamic Communication Topology, Delay Compensation, Multi-Agent Reinforcement Learning.
I Introduction
In recent years, with the rapid deployment of Vehicle-to-Vehicle (V2V) communication facilities and autonomous driving technologies, connected vehicle platoon has become a promising approach for enhancing road safety and traffic efficiency [1]. Through V2V Communication, connected vehicles can form self-organized platoons and maintain a short inter-vehicle distance and coordinated driving behavior, thereby improving road throughput while reducing energy consumption and emissions [2].
In such platoon systems, the effectiveness of cooperative control critically depends on the real-time sharing of vehicle driving state information (e.g., position, velocity, and acceleration, etc.). Consequently, reliable V2V communication is a essential prerequisite for ensuring control performance [3]. However, in real-world traffic environments, channel fading, signal obstruction, and spectral interference may induce substantial communication quality variations, resulting in time-varying delay and packet loss.
These imperfections not only hinder timely and accurate information exchange, as delay leads to outdated data and packet loss results in missing information, but also directly affect the velocity stability and driving comfort of the platoon system [4]. Hence, developing robust platoon control strategies under non-ideal communication conditions has become a pressing research challenge.
To mitigate delay effects, a range of delay compensation based robust control strategies have been proposed. Some studies adopt deterministic or random delay modeling, allowing controllers to compensate for delay-induced performance degradation [5, 6]. More recently, researchers have also explored solutions within the Reinforcement Learning (RL) framework by expanding the agent’s state space to include historical control actions[7, 8], which improves the system’s robustness under delayed communication. Nevertheless, these methods largely rely on simplified delay assumptions and lack explicit sensitivity to the time-varying nature of delay, limiting their scalability to highly dynamic communication environments.
For packet loss, mitigation strategies typically rely on holding previous inputs, interpolating historical data, or fallback modes [9, 10, 11, 12, 13, 14]. Although these methods can maintain basic control performance under mild packet loss, they often fail to guarantee stability and safety in high-loss or highly dynamic environments. Some recent studies have introduced dynamic communication topology mechanisms that adjust weights based on communication quality, aiming to mitigate performance degradation caused by communication disruptions [15, 16]. While this mitigates performance degradation from communication packet loss, these strategies focus solely on communication quality metrics and overlook the need to consider the relevance of communication vehicles in the current dynamic traffic environment when adjusting the communication topology. For example, when a human driver accelerates, they pay attention to the vehicle ahead, while when decelerating, they focus on the vehicle behind[17].
Moreover, most existing studies typically address either communication delay or packet loss in isolation, lacking a unified control framework that considers their coupled impact. Although some traditional optimization-based works have proposed integrated modeling approaches [9, 18, 5, 19, 20], they typically rely on accurate system models and analytical formulations, which limits their adaptability and generalization under real-world uncertainties and unmodeled dynamics. Therefore, it’s necessary to investigate novel intelligent algorithms to solve the adaptability problem in dynamic and complex traffic scenarios.
In contrast, Multi-Agent Reinforcement Learning (MARL), has recently emerged as a promising solution for cooperative platoon control due to its model-free nature, strong generalization capabilities, and support for distributed decision-making [21, 22, 23]. Under assumptions of ideal communication or deterministic delay, previous studies have validated MARL’s efficacy in platoon control [24, 25]. Some methods further enhance MARL by introducing communication-aware designs, that though consider inter-agent information correlations to build dynamic communication topologies [26]. However, these approaches tend to focus on correlation among agents, while overlooking the dynamic of communication quality.
To bridge this gap, this paper proposes a novel Dynamic Communication Topology based Multi-Agent Reinforcement Learning (DCT-MARL) algorithm to enhance the control performance of connected vehicle platoon under non-ideal communication conditions. By dynamically adjusting the communication topology and augmenting the state space, the proposed method improves the robustness and adaptability of vehicle platoon in realistic communication environments. The main contributions are summarized as follows:
  • 1)
    A multi-key gated communication network is proposed, enabling each agent to infer the relevance of communication agents based on real-time observations and historical context. This mechanism dynamically adjusts the communication topology in response to packet loss.
  • 2)
    A state augmentation scheme is designed to mitigate the impact of communication delay on platoon control. Specifically, the agent’s state is extended to include both the history control action and delay information. This augmentation enhances the policy’s sensitivity to time-varying communication dealys and improves the system’s robustness and adaptability under delayed communication.
  • 3)
    A co-simulation environment integrating vehicular control and wireless communication is developed to enable comprehensive evaluation under realistic conditions. Extensive experiments demonstrate that the proposed DCT-MARL algorithm outperforms state-of-the-art methods in terms of string stability and driving comfort under non-ideal communication conditions.
The remainder of this paper is organized as follows: Section II introduced related work on platoon control under non-ideal communication scenarios, as well as MARL-based control approaches. Section III formulates the problem. The details on proposed DCT-MARL based control method is discussed in Section IV. Section V presents the experimental setting and Section VI discusses the experimental results. Finally, the conclusions and future works are given in Section VII.
II Related Work
II-A Platoon Control under Unideal Communication Conditions
Reliable V2V communication is essential for cooperative platoon control, however, packet loss and delays in practical V2V communication may induce performance degradation in the control system[4, 3].
To mitigate the negative impact of communication delay, a range of delay compensation based robust control strategies have been proposed[5, 6, 7, 8, 27]. To address packet loss, traditional approaches employ techniques include Zero-Order Hold (ZOH) to maintain previous control inputs [9], historical data interpolation [10, 11], or fallback modes such as Adaptive Cruise Control (ACC) [12, 13, 14]. Although these methods can maintain basic control feasibility under mild packet loss, they often fail to ensure stability in scenarios with the phenomenon of fluctuating velocities and persistent packet loss. Moreover, predictive strategies may impose substantial computational burdens. To enhance adaptability, some recent studies have introduced dynamic communication topology mechanisms that adjust weights based on communication quality, thereby mitigating the effects of link failures [15, 16]. For example, Liu et al. [15] developed an traditional optimization-based method to adjust weights according to communication conditions, whereas Wang et al. [16] proposed a self-adaptive topology control method based on a Dual-Decision Deterministic Policy Gradient ($D^{3}$PG) algorithm for platoon control, which significantly enhances system resilience under fluctuating communication conditions. However, these strategies primarily focus on communication quality metrics, overlooking the fact that in practical applications, the selection of communication targets should be context-aware and responsive to dynamic traffic conditions.
Furthermore, most existing works consider either delay or packet loss in isolation, and are thus inadequate for scenarios where both coexist. In practice, delay and packet loss often appear simultaneously, imposing compounded challenges. To address this, some recent works have proposed unified models that jointly consider the effects of delay and packet loss[9, 18, 5, 19, 20]. For instance, Vegamoor et al.[18] modeled delay and packet loss simultaneously, and introduced a safety-aware constraint on distance errors between any two vehicles and derived a lower bound on safe time gaps to ensure string stability under unideal communication scenarios. Zhao et al.[5] investigated control strategies under constant time headway distance, modeling non-ideal conditions including limited communication range, random packet loss, and time-varying delay. Their approach employed a distance-based dynamic topology and a robust controller that utilized multi-leader an follower information. Following this, Wang et al.[20] further refined the approach by proposing a CAV following model based on dynamic weight optimization, which adaptively adjusts the importance of incoming information. Their framework incorporated realistic vehicle dynamics and demonstrated strong robustness under complex and adverse communication conditions.
While these control approaches provide valuable theoretical foundations, they often depend on accurate system modeling and incur significant computational overhead. As a result, they are difficult to scale in dynamic and uncertain traffic environments characterized by rapidly changing velocities and nonlinear interactions.
II-B MARL-based Platoon Control
In recent years, MARL has gained significant attention in intelligent transportation systems due to its strengths in distributed decision-making, policy learning, and environmental adaptability[24, 28, 29, 25]. MARL has become a promising paradigm for cooperative control tasks, such as vehicle platoon. For example, Chen et al.[28] evaluated MARL frameworks using both local and global reward structures in platoons of three and five vehicles, and found that decentralized learning with local rewards achieved superior performance in terms of stability and responsiveness compared to centralized approaches.
However, the vast majority of existing MARL-based platoon control studies assume ideal V2V communication, free of delay and packet loss, and thus lack a comprehensive evaluation into the robustness of learned policies under realistic communication impairments. To address this gap, Wang et al. [25] proposed a delay-aware MARL framework with a fixed communication topology, formulated within the partially observable Markov game (POMG), capturing delay-induced uncertainties in platoon systems. Their simulations demonstrated improved platoon stability under time-varying delay. However, existing works still fall short of simultaneously addressing communication delay and packet loss in a unified MARL framework.
On dynamic communication topologies within MARL, researchers have explored inter-agent information correlations to construct adaptive communication topologies. Seminal algorithms including ITGCNet[30], Gated-ACML[31], I2C[26], and NeurComm[32]. In particular, Ding et al. [26] propose I2C, a dynamic communication framework where each agent learns to infer its communication targets from local observations via a prior network trained through causal inference. However, these methods primarily emphasize inter-agent information relevance while neglecting the dynamic variability of the communication environment, limiting their applicability under real-world communication environment.
To overcome these limitations, this paper proposes a DCT-MARL framework that jointly addresses communication delay and packet loss. The framework aims to substantially enhance the robustness and adaptability of connected vehicle platoons in non-ideal communication environment.
III Problem Formulation
III-A Vehicle Dynamics Model
This paper considers a platoon consisting of $N+1$ vehicles, indexed as $0,1,\ldots,n$, where the leading vehicle is indexed as 0, and the following vehicles are indexed from $1$ to $n$. All vehicles communicate through V2V wireless communication to exchange information and coordinate movements. This paper primarily focus on a linear model of a connected automated vehicle (CAV) system, which has been extensively used in vehicle control. The vehicle dynamics are given by [33].
$\begin{gathered}p_{i,t+1}=p_{i,t}+\Delta t*v_{i,t}\\ v_{i,t+1}=v_{i,t}+\Delta t*acc_{i,t}\\ acc_{i,t+1}=\left(1-\frac{\Delta t}{\tau_{i}}\right)acc_{i,t}+\frac{T}{\tau_{i}}*u_{i,t},\end{gathered}$ (1)
where $p_{i,t+1},v_{i,t+1}$ and $acc_{i,t+1}$ represents the position, velocity, and acceleration of vehicle $i$ at time step $t+1$, respectively. The $\Delta t$ is the sampling time. The term $u_{i,t}$denotes the control input (i.e., commanded acceleration), and $\tau_{i}$ is a time constant representing driveline dynamics, describing how the acceleration responds to the control input.
This paper adopt the constant time headway policy (CTHP), meaning that the platoon maintains a constant time gap between vehicles. Each following vehicle $i$ aims to maintain a desired headway distance from its preceding vehicle i1i-1, which is defined as the bumper-to-bumper distance. The ideal inter-vehicle distance between vehicle $i$ and its predecessor i1i-1 is denoted as $d_{r,i,t}$.
$d_{r,i,t}=d_{0}+h*v_{i,t},$ (2)
where $h>0$ is the constant time headway, and $d_{0}$ is the standstill distance that ensures safety when vehicles are stopped. The actual inter-vehicle distance between vehicle $i$ and vehicle i1i-1 is given by $d_{i,t}=p_{i-1,t}-p_{i,t}-l$, where $l$ is the vehicle length. The distance error for vehicle $i$ is then computed as:
$err_{p_{i},t}=d_{i,t}-d_{r,i,t}=p_{j,t}-p_{i,t}-l-d_{0}-h*v_{i,t}.$ (3)
Since all vehicles in the platoon are assumed to be identical, this paper consider $d_{0}$ and $l$ to be uniform across all vehicles and omit them in subsequent derivations.
For any vehicle pair $p_{(}i,j)$ in the communication graph, the tracking error consists of both distance error and velocity error, defined as follows:
$\mathrm{err}_{p_{i,t}}$ $pi,tpj,tμ=i+1jhvμ,t,j>i\displaystyle$= (4)
[…]
where $j<i$ (i.e., vehicle $j$ is ahead of vehicle $i$), the distance error between vehicles $i$ and $j$ is the difference between the actual position minus the velocity of all vehicles from $j+1$ to $i$ multiplied by the sum of $h$.
Each vehicle $i$ is capable of communicating with other vehicle $j$ within its communication range. However, due to potential communication disruptions or packet loss, the communication link may become unavailable at certain time steps. Specifically, if $(i,j)\notin E_{t}$ at certain time steps, vehicle $i$ fails to obtain information from vehicle $j$.
Additionally, due to network congestion, signal attenuation, or interference, communication may experience a delay $\xi_{i,j}$, meaning that vehicle $i$ at time $t$ receives information from vehicle $j$ that corresponds to time $t-\displaystyle\xi_{i,j}$ instead of real-time data.
III-B Control Objectives of Vehicle Platoon
The primary control objective of the connected vehicle platoon is to minimize the deviation of each follower vehicle from its desired relative position and velocity with respect to its preceding vehicles even under non-ideal V2V communication conditions including delay and packet loss. Specifically, the system aims to minimize both the inter-vehicle distance error and the velocity control error, thereby ensuring safe, stable, and efficient following behavior under communication constraints.
Considering the vehicle dynamics described in Eq. 4, the control objective is to minimize the cumulative tracking error over a finite time horizon $T$:
$\min_{u_{i,t}}\sum_{t=0}^{T}\sum_{i=1}^{N}\text{err}_{i,t},$ (5)
where the instantaneous tracking error for vehicle $i$ at time $t$ is defined as:
$\text{err}_{i,t}=\lambda_{p}\|e_{p_{i},t}\|^{2}+\lambda_{v}\|e_{v_{i},t}\|^{2},$ (6)
and $\lambda_{p}$ and $\lambda_{v}$ are weighting coefficients that determine the relative importance of position and velocity tracking errors. These coefficients are typically set to equal values unless a specific prioritization is needed.
In addition, each vehicle’s behavior must satisfy physical and safety-related constraints:
$\begin{gathered}\mathrm{acc}_{\min}\leq\mathrm{acc}_{i,t}\leq\mathrm{acc}_{\max},\end{gathered}$ (7)
where $\mathrm{acc}_{i,t}$ is the actual acceleration of vehicle $i$, and ui,tu{i,t} is the control input (i.e., the commanded acceleration). The bounds $\mathrm{acc}_{\min}$ and $\mathrm{acc}_{\max}$ represent the allowable acceleration range. These constraints ensure safe and comfortable driving and prevent excessive control efforts.
IV DCT-MARL Based Algorithm
In this section, the proposed DCT-MARL algorithm is described, as shown in Figure.1. The platoon control task is then formulated as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP). Furthermore, a dynamic communication topology mechanism is incorporated into the DCT-MARL framework to enhance the robustness of control performance under non-ideal communication conditions.
Figure 1: The architecture of DCT-MARL based Platoon Control algorithm.
The proposed DCT-MARL algorithm is built upon the actor-critic architecture. Each actor consists of a communication network, which is a multi-key gated network based on an inference mechanism. and an action policy network, which determines control actions. Each agent acquires state information by communicating with vehicles within its communication range. Considering potential communication failures, each agent dynamically selects a communication topology at every control cycle based on its local observation, integrating its own state to make control decisions.
This design is well aligned with real-world vehicle platoon systems, where the communication topology is frequently affected by wireless interference, signal blocking, and vehicle mobility. The dynamic topology selection mechanism allows each vehicle to adaptively adjust its communication partners, ensuring that critical information is exchanged even under non-ideal communication conditions.
The centralized critic receives the joint actions and state representations from all agents as input to estimate each agent’s action-value function at each time step. An attention mechanism is utilized to extract relevant information from other agents’ observation-action pairs, enabling each agent to selectively filter critical environmental information. The critic network is trained via temporal difference learning[34] and is only used during training. This paper optimize agent $i$’s policy by minimizing the Kullback–Leibler (KL) divergence $D_{KL}(q_{\theta_{i}}\|p)$, which is given by:
$-D_{\text{KL}}(q_{\theta_{i}}\parallel p)=$ $\sum_{t}\mathbb{E}_{s_{t},u_{t},A_{t}\sim q_{\theta_{i}}}\Big{[}r_{i}(s_{t},u_{t})+$ (8)
$H\big{(}\pi_{i}(u_{i}\mid o_{i}(s,A))\cdot\rho(A_{t}\mid s_{t})\big{)}\Big{]}.$
Beyond the reward term, the objective introduces an additional conditional entropy term for the joint distribution $\pi_{i}(a_{i}|o(s_{t},A_{t}))\rho(A_{t}|s_{t})$, which encourages exploration for agent $i$ and enhances graph policy $\rho$. The objective function for agent $i$ is now redefined as:
$J_{\pi_{i}}(\theta_{i})=\mathbb{E}_{q_{\theta_{i}}}\left[Q(s,u)+H\left(\pi_{i}(u_{i}|o_{i}(s,A))\rho(A|s)\right)\right].$ (9)
The policy $\pi_{i}$ is updated by maximizing $J_{\pi_{i}}$, and its gradient is:
$\nabla_{\theta_{i}}J(\theta_{i})=\mathbb{E}_{q_{\theta_{i}}}\Big{[}$ $\nabla_{\theta_{i}}\log\pi_{i}(u_{i}\mid o_{i}(s,A))\cdot$
$\left(-\alpha\log\pi_{\theta_{i}}(u_{i}\mid o_{i}(s,A))+Q_{i}(s,u)\right)\Big{]}.$ (10)
The loss function for $Q_{i}$ is defined as:
$L_{Q}(\phi)=\mathbb{E}_{(s_{t},u_{t})\sim D}\left[(y_{t}^{i}-Q_{i}(s_{t},u_{t};\phi))^{2}\right],$ (11)
where
$y_{t}^{i}=r_{i}(s_{t},u_{t})+Q_{i}(s_{t+1},u_{t+1}^{i},u_{t+1}^{-i};\phi).$ (12)
During training, we follow these iterative procedures:
  • First, fix $\pi_{i}(u_{i}|s,A)$ and $Q_{i}(s,u)$ to update $\rho(A)$.
  • Then, improve $\pi_{i}(u_{i}|o(s,A))$ and $Q_{i}(s,u,A)$ using the generated trajectories given $A$.
The prior network, parameterized by $\theta_{b_{i}}$, is trained as a binary classifier using the training dataset $S$ and is updated with the following loss function:
$\mathcal{L}(\theta_{b_{i}})=\mathbb{E}_{(o_{i},d_{j}),\mathcal{I}_{i}^{j}\sim\mathcal{S}}\Big{[}$ $-(1-y_{i}^{j})\log\left(1-b_{i}(o_{i},d_{j})\right)$
$+y_{i}^{j}\log\left(b_{i}(o_{i},d_{j})\right)\Big{]},$ (13)
where, $y_{i}=\begin{cases}1,&\text{if }P_{\omega}(y_{i}\mid o_{i})\geq\delta\\ 0,&\text{otherwise},\end{cases}$ and $\delta$ is a hyperparameter. Notably, if a particular agent plays a critical role in shaping agent $i$’s state estimation during centralized training, its information should remain accessible during decentralized execution. To this end, a dynamic communication topology based on communication target selection is proposed and shared across both training and execution phases, enabling robust and adaptive coordination.
The proposed mechanism employs a multi-key gating structure to dynamically select communication targets, accommodating practical V2V constraints such as delay and packet loss. By adjusting the topology in response to traffic context, it ensures stable coordination and safety performance of platoon systems under complex and dynamic environments.
IV-A Dec-POMDP Formulation
A Dec-POMDP is formally defined as a tuple $\langle N,S,U,P,R,O,\gamma\rangle$, where $N={1,\dots,n}$ denotes a finite set of $n$ agents interacting in a shared environment. The environment is characterized by a global state space $S$, and each agent selects actions from a joint action space $U$. The environment dynamics are governed by the state transition function $P:S\times U\rightarrow S$, which defines the probability distribution over subsequent states given the current state and joint actions. Each agent receives observations from an observation space $O$, which may provide only partial and possibly noisy information about the true state. The reward function $R:S\times U\rightarrow\mathbb{R}$ maps the current state and joint actions to a scalar reward shared among all agents. Finally, $\gamma\in[0,1)$ is the discount factor that determines the present value of future rewards.
At each time step $t$, each agent $i\in N$ receives an observation from the state and follows an individual policy $\pi_{i}$, where $\pi_{i}$ depends on the agent’s observation-action history. The joint action of all agents leads to the next state, and the team receives a global reward. Each agent aims to maximize the accumulated reward.
State Space $S$: To mitigate the performance degradation caused by communication delay, this paper extend the definition of the agent’s state space by incorporating both historical control information and delay descriptors. Specifically, at time step $t$, the state of vehicle $i$ is defined as:
$S_{i,t}=\left\{p_{i,t},v_{i,t},u_{i,t},u_{t-1},\xi_{i,t}\right\},$ (14)
where $p_{i,t}$, $v_{i,t}$, and $u_{i,t}$ denote the vehicle’s current position, velocity, and acceleration, respectively; $u_{t-1}$ represents the most recent control input, and $\xi_{i,t}$ denotes the communication delay associated with the latest received information.
Although this formulation deviates from the conventional state definition, it is a rational extension under communication-impaired environments, where both the delayed control input and latency should be explicitly considered in the state representation. In connected vehicle systems, time-varying wireless communication often causes neighboring information to arrive with variable delay. If the control policy relies solely on such delayed observations, the Markov property may be compromised, thereby impairing decision-making performance.
By incorporating the previous control action $u_{t-1}$, the agent can better infer the underlying system dynamics during periods of delayed or missing observations. Additionally, explicitly modeling the delay $\xi_{i,t}$ enables the policy to assess the reliability of received data and adjust its decisions accordingly. Upon receiving information from neighboring agents, the agent estimates the communication delay using the attached timestamps and incorporates this value into its current state. This augmented state representation preserves the Markovian assumption under delayed observations and significantly enhances decision-making robustness in non-ideal communication environments.
Observation Space $O$: Agents use communication to obtain information beyond their direct observations to assist decision-making. Each agent’s local observation $o_{i}\in S_{i}$ is a partial description of the environment state. At time step $t$, each agent $j$ in the platoon transmits its driving information $X_{j,t}={p_{j,t},v_{j,t},u_{j,t},t_{j,t}}$, neighboring agent $i$ obtains driving information from multiple vehicles within its V2V communication range, forming its partial observation $o_{i}$ and determining which agents are within communication range. This solves the problem of possible communication failures. For example, if agent l fails within the current control interval k, agent i cannot observe agent l’s information through V2V communication, and can receive the status information of other related vehicles for control output. It avoids the problem that the amount of states received may be insufficient due to communication failures. The global state $O$ is then defined as the combination of all agents’ local observations: $O=\{o_{1},o_{2},\dots,o_{n}\}.$ Similarly, the global state $S$ is defined as: $S=\{s_{1},s_{2},\dots,s_{n}\}.$
Action Space $U$: Each agent $i$ follows a decentralized policy $\pi_{i}:O_{i}\times S_{i}\to[0,1],$ to determine the control output at time step $t$. Each PC agent $i$ selects a control input as its PC action at time step $t$.
Reward Function $R$: The reward function $r_{i,t}$ is critical for training RL agents to exhibit the desired behavior. Our objective is to train agents to ensure platoon stability and comfort under imperfect communication and dynamic traffic conditions. The reward assigned to agent $i$ at each time step $t$ is designed as follows:
$r_{i,t}=-\big{\{}$ $\omega_{1}\left(\text{err}_{p_{i},t}\right)+\omega_{2}\left(\text{err}_{v_{i},t}\right)+\omega_{3}\left(\text{jerk}_{i,t}\right)$ (15)
$+\omega_{4}\cdot\operatorname{ReLU}\left(d_{\text{safe}}-d_{i,t}\right)\big{\}},$ (16)
where $w_{1}$, $w_{2}$, $w_{3}$, and $w_{4}$ are the weighting coefficients. In this equation, the first two terms penalize deviations from ideal position and velocity, encouraging agents to achieve these goals closely and guaranteeing the stability of the platoon. The third term is to improve driving comfort to minimize sudden acceleration, thereby providing a smoother and more comfortable ride for passengers. Where jerk is the rate of change of acceleration, which is given by
$\text{jerk}_{i,t}=\frac{acc_{i}(t)-acc_{i}(t-1)}{\Delta t}.$ (17)
The fourth item acts as a safety constraint, severely penalizing agents if the distance between vehicles is less than the safe distance, which is essential to prevent collisions and ensure the safety of the vehicle queue. $d_{i,k}$ is the actual distance between vehicle $i$ and the vehicle in front, and $d_{\text{safe}}$ is the preset minimum safe distance. $\operatorname{ReLU}(\mathrm{x})=\max(0,x)$ used to ensure that a safety item generates a penalty only if the actual distance is below the safety threshold.
In addition, considering the communication delay problem, each agent can only access the lagged observation information when making a decision, and cannot directly obtain the current true state. Therefore, we modify the reward function based on conditional expectation to adapt it to the optimal learning in the communication constrained environment. Specifically, for each agent $i$, in the extended state $S_{i,t}$, the reward function of delay is defined as follows:
$\tilde{R}_{i,t}\left(S_{i,t},u_{i,k}^{\mathrm{CL}}\right)=\mathbb{E}_{x_{i,t}}\left[R_{i,t}\left(x_{i,t},u_{i,t}\right)\mid S_{i,t}\right].$ (18)
This modification ensures that even in a communication-constrained environment, agents can still make reasonable decisions based on past observations and actions while maximizing their long-term rewards.
Therefore, each agent determines its action based on its own observations and the messages received in a distributed manner. This process is repeated for each time step under different observations. This work aims to achieve comfort and safety control of platoon. The ultimate goal of the RL model is:
$\max\left(J_{\pi}(s(0))=\lim_{M\rightarrow\infty}\mathbb{E}\left[\sum_{t=0}^{M-1}\gamma^{k}r_{i}(s(t),\pi(s(t)))\right]\right),$ (19)
where $J_{\pi}(s(0)$ is the expected cost with initial condition $s(0)$ and control policy $\pi$, $\gamma$ is the discounted factor, $M$ is the number of steps; $E(\cdot)$ is the mathematical expectation. The update strategy of the proposed DCT-MARL is given in Algorithm 1.
Algorithm 1 DCT‑MARL: Dynamic Communication Topology‑based Multi‑Agent RL.
Ensure: Action Policy $\pi_{i}$, and graph reasoning policy $\rho$
1: Initialize parameters $\theta_{i}$, $\phi_{i}$ for each agent and $\phi$ for graph reasoning;
2: Assign target parameters of Q-function: $\phi_{i}\leftarrow\phi_{i}$, and target policy parameter: $\theta_{i}\leftarrow\theta_{i}$
3: $D\leftarrow$ empty replay buffer
4: while not terminated do
5:     Sample $A\sim\rho_{\phi}(A|s)$;
6:     Given current state $s$, compute $o_{i}(s,A)$ for all agents and $x$;
7:     Sample $u_{i}\sim\pi_{i}(\cdots|o_{i}(s,A))$;
8:     Take the joint action $\{u_{i}\}_{i\in N}$ and observe own reward $r_{i}$ and new state $s^{\prime}$:
    $D\leftarrow D\cup\{(s,\{u_{i}\}_{i\in N},A,\{r_{i}\}_{i\in N},s^{\prime})\}$;
9:     for each agent do
10:      Update $\phi_{i}$ according to $\nabla_{\phi}J_{Q}(\phi_{i})$;
11:      Update $\theta_{i}$ according to $\nabla_{\theta_{i}}J_{\pi}(\theta_{i})$;
12:      Update $\phi$ according to $\frac{1}{n}\sum_{i=1}^{n}\nabla_{\phi}\eta_{i}(\rho_{\phi})$;
13:      Reset target Q-function: $\phi_{i}=\beta\phi_{i}+(1-\beta)\phi_{i}$;
14:     end for
15: end while
IV-B Dec-POMDP with Dynamic Directed Graph
The vehicles in the platoon exchange their state information (such as position, velocity, and acceleration) through a V2V wireless communication network. Each vehicle has its own communication network, which can be modeled as a directed weighted graph $G=(T,V,E)$, where $t\in T$ is the set of the trajectory in the time dimension, $V$ represents the set of vehicles. $E\subseteq V\times V$ represents the set of communication links, where $(i,j)\in E$ indicates that vehicle $i$ can send information to vehicle $j$, representing the communication connection between vehicles. The adjacency trajectory matrices for the dynamic directed graph and subgraphs are denoted by $\mathbf{A}$. When a node can access the observations of other nodes, it can approximate the global state.
Based on Dec-POMDP, we propose a communication topology using dynamic directed graphs. The structural properties of the graph are represented by the adjacency trajectory matrix $A$, which is denoted as $\mathbf{A}\in\mathbb{B}^{l\times n\times n}$, where B indicates the Boolean matrix, l denotes the length of the trajectory and n indicates the number of agents. If $\lceil A^{ij}_{t}\rceil=0$, then no communication occurred from agent $i$ to agent $j$ at time $t$. If $\lceil A^{ij}_{t}\rceil=1$, then communication transpired from agent $i$ to agent $j$ at time $t$.
IV-C Multi-Key Gating Network Based Dynamic Communication
During communication, an agent intuitively prefers to interact with those agents that significantly influence its decision-making strategy. Specifically, assuming agent $i$ can receive information from agent $j$, it first inputs its local observation $o_{i}$ along with the ID $d_{j}$ of agent $j$ into the model, which then outputs a confidence score indicating the likelihood of communication with agent $j$. To this end, this paper employs a feedforward neural network to implement causal inference, leveraging the KL divergence to quantify the causal confidence between agent $i$ and other agents. The formulation is given as follows:
$D_{\text{KL}}(P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})\|P(a_{i},a_{j}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})),$ (20)
where $P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})$ and $P(a_{i},a_{j}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})$ denote two conditional probability distributions over the action space of agent $i$, and $\boldsymbol{o}$ represents the joint observations of all agents. Here, $\boldsymbol{a}_{-i}$ refers to the joint actions of all agents excluding agent $i$, and similarly, $\boldsymbol{a}_{-ij}$ denotes the joint actions of all agents excluding both agents $i$ and $j$.
The latter probability distribution does not depend on agent $j$’s actions, meaning that when making decisions, agent $i$ ignores the actions of agent $j$. Note that the conditional actions of other agents $a_{-i}$ are sampled from their current policies. The magnitude of KL divergence quantifies how agent $i$ would adjust its own policy if it considered agent $j$’s actions, thus indicating the degree of correlation between agent $j$’s policy and agent $i$’s policy.
The two probability distributions are computed using the joint action-value function, where $P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})$ is derived from the softmax distribution over agent $i$’s actions:
$P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})=\frac{\exp(Q(a_{i},\boldsymbol{a}_{-ij},\boldsymbol{o})/\lambda)}{\sum_{a_{i}^{\prime}}\exp(Q(a_{i}^{\prime},\boldsymbol{a}_{-ij},\boldsymbol{o})/\lambda)},$ (21)
where $Q(\boldsymbol{a},\boldsymbol{o})$ is the joint action-value function, and $\lambda\in\mathbb{R}^{+}$ is the temperature parameter. The marginal distribution $P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})$ is computed as follows:
$P(a_{i}\mid\boldsymbol{a}_{-ij},\boldsymbol{o})=\sum_{a_{j}}P(a_{i},a_{j}\mid\boldsymbol{a}_{-ij},\boldsymbol{o}).$ (22)
Then, a multi-key gating network is used to select $m$ communication targets, where the Gumbel-Softmax activation function determines whether an agent should communicate or not. The computation is given by:
$key_{k}^{i}=\text{ gumbel }-\operatorname{softmax}\left(\mathrm{W}_{\mathrm{v}}\tanh\left(\mathrm{~W}_{\mathrm{q}}\mathrm{o}_{\mathrm{k}-1}^{\mathrm{i}}+\mathrm{W}_{\mathrm{k}}\mathrm{o}_{\mathrm{k}-1}^{\mathrm{i}}\right)\right),$ (23)
where Gumbel-Softmax [35, 36]is an activation function with two output dimensions (communicate or not). The weight matrices $\mathbf{W}_{v}$, $\mathbf{W}_{q}$, and $\mathbf{W}_{k}$ determine the communication importance, considering the dynamically changing communication links in a vehicle platoon scenario due to traffic conditions.
Finally, the adjacency matrix $\mathbf{A}$ is generated by summing over all keys:
$A_{i,k}=\left\lceil\max\left(key_{k}^{i}\mid j\right)\right\rceil,$ (24)
where $key_{k}^{i}\mid j$ represents the total number of keys obtained by agent $i$ in control cycle $k$, and $\lceil\cdot\rceil$ denotes the ceiling function. This ensures that the final communication topology is dynamically adjusted rather than fixed. Additionally, the number of keys ensures that each agent only communicates with a limited number of selected agents, thereby maintaining a stable input dimension.
To address the lack of interpretability in purely data-driven methods, this paper introduce a dynamic weighting module based on safety distance, which assigns higher weights to closer vehicles to reflect their critical role in collision avoidance. Specifically, the weight of vehicle $j$ is defined as:
$\omega_{j,k}=\frac{\frac{1}{2^{|i-j|}}}{\sum_{l\in N_{\mathrm{i},\mathrm{k}}}\frac{1}{2^{|\mathrm{i}-\mathrm{j}|}}},$ (25)
where $d_{ij}$ represents the distance between vehicle $i$ and vehicle $j$, and $\epsilon$ is a small positive constant to avoid division by zero. Finally, all received messages are fed into the message encoder network $e_{i}(\cdot)$ to generate encoded messages $c_{i}$, and the policy network produces the action distribution $\pi_{i}(a_{i}|c_{i},o_{i})$.
V Experimental Setting and Metrics
V-A Experimental Setup
To validate the effectiveness of the proposed control strategy, this paper further incorporates empirical trajectory data from the NGSIM dataset[37], a widely used benchmark in vehicle platoon control research that includes congested traffic scenarios. This enables the evaluation of the controller’s ability to handle stop-and-go traffic oscillations commonly observed in real-world platoon environments. Specifically, the driving profile of the leading vehicle is extracted from real-world trajectory data collected on the eastbound I-80 highway in the San Francisco Bay Area. The use of realistic driving patterns contributes to more effective policy learning for connected vehicles platoon operating under complex and dynamic conditions.
In the communication simulation, the wireless channel model incorporates both small-scale and large-scale fading effects. Small-scale fading is modeled as Rayleigh fading, capturing rapid signal fluctuations due to multipath propagation. Large-scale fading includes path loss (PL) and shadow fading (SF).
Path loss is modeled using the WINNER+ B1 model recommended by 3GPP [38], which defines as:
$\mathrm{PL}(d)=A_{0}+10n\log_{10}\left(\frac{d}{d_{0}}\right)+X_{\sigma},\quad\text{(dB)}$ (26)
where $d_{0}$ is the reference distance; $A_{0}$ is the intercept representing the path loss at the reference distance $d_{0}$; $n$ is the path loss exponent, which varies for different traffic scenarios and is set to $n=2.15$ for highway V2V scenarios; $X_{\sigma}$ represents shadow fading, modeled as a Gaussian random variable with zero mean and variance $\sigma^{2}$, caused by obstacles such as trees, buildings, and vehicles. Formally, $X_{\sigma}\sim\mathcal{N}(0,\sigma^{2})$, where the standard deviation is $\sigma=3$ dB.
The simulation environment models the communication scenario and control algorithms using Python, and integrates with SUMO for co-simulation and optimization via the TraCI interface. The simulation time step and duration are synchronized with those of SUMO. A summary of all simulation parameters is presented in Table I.
TABLE I: Simulation Parameters for Experimental Setup.
Parameters Value
Transmission Power 23 dBm
Number of Subchannels 4
Subchannel Resource Blocks (RBs) 10
Channel Bandwidth 10 MHz (50 RBs)
Packet Size 190 bytes
Message Transmission Interval 100 ms
Simulation Duration 60 s
Simulation Time Step 0.1 s
In the reward function of the DCT-MARL algorithm (Equation 15), the weighting coefficients $\omega_{1}$, $\omega_{2}$, $\omega_{3}$, and $\omega_{4}$ are set to 1.0, 1.0, 0.1, and 10.0, respectively, to emphasize the penalty for unsafe following distances. The algorithm adopts a centralized training and decentralized execution framework. Both the centralized critic and policy networks are implemented with three fully connected layers, while the prior network consists of two fully connected layers. Each hidden layer contains 128 neurons. The activation function used in all networks is the Rectified Linear Unit (ReLU), defined as $\operatorname{ReLU}(x)=\max(0,x)$.
During training, each batch consists of 60 time steps, and the maximum episode length is set to 600 steps. The learning rate for both the actor and critic networks is set to $1\times 10^{-3}$. All hyperparameters were determined through grid search.
V-B Evaluation Metrics
To comprehensively evaluate the performance of the proposed platoon control strategy, this paper adopted two categories of metrics: platoon stability and driving comfort. Platoon stability reflects the system’s ability to suppress inter-vehicle disturbances and avoid amplification of oscillations downstream. Following[21], we define the stability index $S_{i}$ for each vehicle as the time-averaged weighted squared deviation in headway and relative velocity:
$S_{i}=\frac{1}{N_{s}}\sum_{t=1}^{N_{s}}s_{i,cf}(t)Cs_{i,cf}^{T}(t),$ (27)
where $s_{i,cf}(t)=\left[\Delta h_{i,i-1}(t),\Delta v_{i,i-1}(t)\right]$ denotes the deviation vector between vehicle $i$ and its preceding vehicle at time $t$, $C$ is a weighting matrix, and $N_{s}$ is the total number of simulation steps. Lower values of $S_{i}$ correspond to better string stability, reflecting reduced propagation of distance and velocity errors across the platoon.
Driving comfort is assessed based on the smoothness of acceleration changes. The normalized comfort score $G_{i}^{\prime}$ is defined as:
$G_{i}^{\prime}=1-\frac{1}{N_{s}}\sum_{t=1}^{N_{s}}\left(\frac{acc_{i}(t)-acc_{i}(t-1)}{jerk_{\text{max}}\cdot\Delta t}\right)^{2},$ (28)
where $jerk_{\text{max}}$ represents the maximum allowable change rate of acceleration used for normalization, which is set to 1.5 m/s3 according to SAE comfort thresholds [39]. Higher values of $G_{i}^{\prime}$ correspond to smoother acceleration transitions and thus improved driving comfort.
VI Performance Results and Analysis
VI-A Control Performance against Existing Platoon Control Strategies
To verify the effectiveness and robustness of the proposed strategy, this paper conducts a comprehensive comparison between the proposed DCT-MARL method and other state-of-the-art baselines: the traditional optimization-based Dynamic Weights Optimization based Control (DWOC) strategy [20], which considers both communication delay and packet losse by dynamically compensating for front-vehicle information through link-quality-aware weighting; the DRL-based $D^{3}$PG method [16] proposed a self-adaptive topology control method that significantly improves resilience under fluctuating link conditions. To ensure a fair comparison, the state space setting of $D^{3}$PG is aligned with our proposed method, including the incorporation of delay compensation mechanism.
Fig. 2 illustrates the temporal responses of vehicle velocity and acceleration under an oscillatory driving scenario. Across all methods, a noticeable attenuation of acceleration fluctuations is observed along the platoon, indicating partial suppression of error propagation. However, notable discrepancies persist in the response to abrupt velocity variations and in the smooth attenuation of acceleration during convergence.
(a) DWOC Based Platoon Control
(b) $D^{3}PG$ Based Platoon Control
(c) DCT-MARL Based Platoon Control
Figure 2: Performance Results of Vehicle Platoon Under Different Strategies.(a) DWOC (b) $D^{3}PG$ (c) DCT-MARL
The DWOC strategy (Fig. 2 (a)) achieves control stability by adjusting the weights assigned to front vehicle information. This approach provides a certain degree of robustness against delay and packet loss. However, sharp acceleration changes in acceleration are observed between 15s and 25s, indicating challenges in maintaining smooth transitions during sudden maneuvers of the lead vehicle. In contrast, $D^{3}PG$ (Fig. 2 (b)) improves inter-agent coordination by selecting communication neighbors using attention-based mechanisms. Nonetheless, due to its single-agent DRL design lacking full MARL coordination, downstream vehicles show delayed reactions to upstream disturbances. The proposed DCT-MARL method (Fig. 2 (c)) consistently delivers the most stable and coordinated control responses, featuring minimal velocity fluctuations and enhanced inter-vehicle coherence, thereby demonstrating superior robustness under non-ideal communication conditions.
As shown in Table II, DWOC and $D^{3}$PG maintains consistent performance across the platoon; however, DWOC disturbances tend to amplify toward the rear of the platoon, indicating limited suppression capability. Although $D^{3}$PG alleviates this issue via communication-aware strategies, intermittent oscillations still occur under uncertain communication conditions. In contrast, DCT-MARL achieves the lowest string stability metrics across all vehicles, particularly in $CAV_{3}$ to $CAV_{5}$, demonstrating its superior ability to suppress disturbance propagation. Moreover, DCT-MARL yields the highest overall driving comfort scores, maintaining low acceleration variance even for the last vehicle in the platoon, which is typically most susceptible to disturbance, thereby ensuring an improved passenger experience.
TABLE II: Quantified Metrics for Each CAV in the Platoon Controlled by Different Policys.
$CAV_{1}$ $CAV_{2}$ $CAV_{3}$ $CAV_{4}$ $CAV_{5}$
String Stability DWOC 0.109 0.177 0.266 0.295 0.312
$D^{3}PG$ 0.097 0.141 0.154 0.258 0.221
DCT-MARL 0.086 0.097 0.088 0.112 0.103
Driving Comfort DWOC 0.512 0.404 0.358 0.329 0.302
$D^{3}PG$ 0.643 0.616 0.592 0.571 0.548
DCT-MARL 0.827 0.816 0.797 0.771 0.775
To illustrate this more clearly, Fig. 3 presents the performance comparison of various control strategies. The DCT-MARL curve maintains a relatively flat comfort trajectory, while all other baselines show progressively deteriorating performance over time. This highlights DCT-MARL’s robustness and resilience under dynamic and uncertain communication conditions. By jointly considering topology adaptation, delay compensation, and cooperative policy learning, the DCT-MARL method achieves superior multi-objective optimization, balancing safety, comfort, and stability more effectively than DWOC, and $D^{3}PG$. In summary, the DCT-MARL model has been proven to be effective in improving road traffic capacity.
Figure 3: Quantified Metrics of Dfferent Models.
VI-B Control Performance against Existing MARL Algorithms
To further assess the advantages of the proposed DCT-MARL framework, we compare it against two representative MARL algorithms: Delay-Compensated MARL (DEC-MARL)[25], which explicitly addresses communication delay by extending the state space to incorporate historical control outputs under a fixed communication topology. Learning-based MARL (LEB-MARL)[26], which introduces a framework to learn communication priors through causal inference and regularizes agent policies to enhance message utilization.
To assess learning efficiency, the average episodic return is depicted in Fig. 4, which compares the performance of the proposed DCT-MARL algorithm against other MARL methods. As training progresses, DCT-MARL consistently achieves higher rewards, indicating more effective policy optimization. In contrast, DEC-MARL and LEB-MARL exhibit slower convergence and lower final reward values, reflecting reduced learning efficiency in dynamic communication environments. These results demonstrate that DCT-MARL outperforms the baselines in both convergence speed and reward performance, highlighting its superior adaptability and learning capability.
Figure 4: Average Episode Reward vs. Training Steps of Various Methods.
Table III compares the control performance of the three methods in terms of string stability and driving comfort for each vehicle within the platoon. While DEC-MARL enhances system stability to a certain extent under non-ideal communication conditions, the reliance on a fixed communication topology and absence of dynamic adjustment mechanisms cause significant performance degradation under high packet loss rates and dynamic communication structures. This results in greater spacing fluctuations among following vehicles and a noticeable decline in driving comfort. LEB-MARL incorporates agent correlation and demonstrates certain advantages in control accuracy. However, it fails to account for the stochastic nature of communication links, limiting robustness under packet loss and topological disturbances and leading to more pronounced oscillations in string stability. In contrast, DCT-MARL consistently delivers more robust and stable control outcomes across all vehicles in the platoon.
TABLE III: Quantified Metrics for Each CAV in the Platoon Controlled by FCT-MARL and DCT-MARL.
$CAV_{1}$ $CAV_{2}$ $CAV_{3}$ $CAV_{4}$ $CAV_{5}$
String Stability DEC-MARL 0.114 0.138 0.167 0.248 0.279
LEB-MARL 0.103 0.125 0.134 0.197 0.221
DCT-MARL 0.086 0.097 0.088 0.112 0.103
Driving Comfort DEC-MARL 0.681 0.628 0.538 0.516 0.505
LEB-MARL 0.739 0.712 0.6358 0.673 0.629
DCT-MARL 0.827 0.816 0.797 0.771 0.775
Besides training performance, the runtime efficiency of the proposed DCT-MARL is also assessed during the testing phase. On a local testing platform equipped with a 2.0 GHz quad-core Intel Core i5 processor and 16 GB of 3733 MHz LPDDR4X memory, executing a complete 600-step simulation required approximately 3.7 seconds. This corresponds to an average inference time of around 6 ms per control step, which is significantly below the 100ms control horizon. These results demonstrate that the proposed method not only achieves superior control performance but also possesses high inference efficiency, making it well-suited for real-time applications or deployment scenarios with limited computational resources.
VI-C Eiffectiveness of Dynamic Communictaion Topology
To further examine the necessity and rationality of the dynamic communication topology mechanism, Fig. 5 illustrates the communication heatmap among six platoon memebers during the testing phase.
Figure 5: Communication Heatmap of CAV Platoon with Dynamic Topology.
In the heatmap, the color intensity indicates the frequency of message exchanges between agents. The vertical axis represents the message sender, while the horizontal axis corresponds to the receiver. In this visualization, a darker color indicates a higher likelihood of information exchange, reflecting stronger communication links between specific vehicle pairs. This allows us to intuitively assess how communication patterns evolve under the dynamic topology.
Notably, as the leading vehicle operates as a non-controllable preset agent, it is only possible to observe whether other vehicles receive messages from it, without direct access to the messages received by the leader itself. Accordingly, the corresponding cells representing the leader as a receiver are shaded in light gray to indicate the absence of observable data.
It can be observed that while front-vehicle information receives predominant attention in the control process, rear-vehicle data also demonstrates significant utilization. This observation is consistent with the findings of[5], highlighting the latent value of rear-vehicle states in control decision-making. Compared to traditional fixed topologies, the dynamic topology can adaptively adjust communication links in real-time based on control needs and information value. These results demonstrate the meaningfulness and effectiveness of incorporating dynamic communication topology under non-ideal communication conditions.
VII Conclusions and Future Works
This paper addresses connected vehicle platoon control under non-ideal communication conditions characterized by time-varying delays and packet loss. To tackle these challenges, DCT-MARL is proposed. Specifically, a multi-key gated communication mechanism is introduced, which dynamically adjusts the communication topology based on the correlations among agents and their current communication status to effectively mitigate packet loss. Furthermore, a state augmentation scheme is designed to incorporate historical control actions and communication delay information into each agent’s state representation, ensuring the preservation of the Markov property despite delayed observations. This enhancement further improves the robustness of the system in realistic V2V communication environments. Experimental results demonstrate that the proposed approach significantly improves platoon stability and comfort, outperforming existing state-of-the-art baselines across multiple key performance metrics. Future research will further investigate the mechanism of non-ideal communication and explore the joint optimization of communication and control, enabling more adaptive and resource-efficient platoon coordination.
Acknowledgments
This work was supported by Beijing Natural Science Foundation under Grant L202018, the National Natural Science Foundation of China under Grant 61931005, the Key Laboratory of Internet of Vehicle Technical Innovation and Testing (CAICT), Ministry of Industry and Information Technology under Grant No. KL-2023-001.